OpenStudy (anonymous):

Let f(X)=cos2x. Find a quadratic polynomial px so that p0=f0, p'0=f'0 and p''0=f''0. px=

OpenStudy (ash2326):

Let's find f(0)= cos 0=1= p(0) P(x) must have a constant term let it be 1 Let's find f'(x) \[f'(x)= -2\sin 2x\] x=0 \[f'(x)=0\] p'(0)=0 Let's find f''(x) \[ f''(x)= -4 cos 2x\] \[ f''(0)=-4\] so p''(0)=-4 p"(x)= -4 We can p'(x)= -4x+c p(x)=-2x^2+c1x+c2 p(0)=1 c2=1 p'(0)=0 c2=0 p''(0)=-4 so \[p(x)= 1-2x^2\]

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

ddx27 did you understand?

OpenStudy (anonymous):

yea i just had a problem with findin f''(0)

OpenStudy (anonymous):

\[f(0) = 1 \] \[f \prime(x) = -2\sin2x \] \[f \prime(0) = 0\] \[f \prime \prime(x) = -4 \cos 2x\] \[f \prime \prime(0) = -4\] \[p(x)=ax ^{2}+bx+c\] \[p(0) = c\] \[p \prime(x)=2ax+b\] \[p \prime(0) = b\] \[p \prime \prime(x) = 2a\] \[p \prime \prime(0) = 2a\] \[Equating\] \[c = 1\] \[a = -2\] \[b = 0\] \[p(x) = -2x ^{2} + 1\]